# EcoServe: Designing Carbon-Aware AI Inference Systems

### v1

@@ III-A2 Memory and Storage
Recent efforts from industry demonstrate memory and storage dominate the embodied carbon of Microsoft’s data center hardware [69]. To quantify the memory and storage embodied carbon, we consider both memory technologies and capacities. For storage, we estimate SSD emissions based on the Dell R740 life cycle analysis [10], yielding an average estimate of 0.110 kg CO2e per GB capacity; this is a conservative estimate compared to estimates of 0.160 kg CO2e per GB from academic surveys [64]. For memory, we extend existing ACT and SCARIF models by scaling wafer-scale DRAM data for different technologies from TechInsights [35, 66, 3] by publicly available DRAM bit-densities per technology from memory vendors (e.g., SK Hynix, Micron, Samsung) as shown in Figure 2. The trends show newer technology nodes, given higher bit-densities, exhibit lower embodied carbon per GB capacity. For the purposes of this work we estimate the DDR4/LPDDR5, GDDR6, and HBM2, HBM3e embodied carbon footprint as 0.29, 0.36, 0.28, 0.24 kg CO2e per GB respectively.
III-A3 Mainboard Printed Wiring Board (PCB)
To estimate the embodied emissions of the mainboard printed wiring board (PWB) we use the Dell R740 LCA [10]. Given the PWB area of the Dell R740, 1925 square centimeters (176 kg CO2), is significantly higher than modern server-class host systems and GPU’s, we scale the PWB area for specific systems. Thus, we estimate the embodied carbon of PWB to be: Total PCB $\text{Total PCB CF}=0.048\times Area_{PCB}$.
@@ III-A3 Mainboard Printed Wiring Board (PCB)
Figure 3: Trends in embodied carbon, power, and cloud cost for different generations of GPUs. As GPU performance increases (left to right), power consumption, cost and embodied carbon exhibit distinct trends.  [8, 9, 1].
III-B Characterizing Total Carbon
@@
Figure 3 analyzes the embodied carbon estimates of various NVIDIA GPUs, breaking it down into key components such as memory, printed circuit board (PCB), power delivery network (PDN), cooling systems, and system-on-chip (SoC). Figure 3 illustrates these components using stacked bars to represent embodied carbon, with the thermal design power (TDP) plotted on the right axis. We observe a clear trend: with newer GPU generations, computational performance, memory bandwidth, and energy efficiency improve. However, the embodied carbon also rises with increasing computational performance, advanced thermal solutions required, and improved memory capacity. For instance, compared to an NVIDIA H100, and NVIDIA L4 incurs 3$\times$ lower embodied carbon.
@@
Figure 4: Embodied carbon breakdown of full inference systems available in cloud offerings from Azure and LambdaLabs [9, 1], varying the number and type of GPU. Host-processing systems account for over half of the embodied carbon in AI systems, owing largely to memory, storage, and mainboard overheads.
Figure 4 shows the embodied carbon breakown of end-to-end inference systems including both GPUs and host systems. The analysis highlights that the motherboard, DRAM, and storage constitute a significant portion of the embodied carbon impact. However, for AI inference workloads, host systems are frequently underutilized compared to GPUs, leading to a disproportionate contribution of their embodied carbon footprint relative to their actual usage during operations. This observation motivates exploring strategies such as resource sharing on the host (e.g., model multiplexing), improving utilization through CPU-based inference, extending the lifetime of host system components, and reducing the memory subsystem’s footprint (see Section IV).
@@
Figure 5 compares the embodied carbon versus the operational carbon per second across datacenters with different power sources. The host and GPU embodied carbon remained fixed, and the operational carbon footprint varies with the carbon intensity of energy sources. We show that for regions with high carbon intensity (Europe and US-East), the GPU operational carbon dominates the total footprint. As we move to cleaner energy sources, the operational emissions drop, leaving the embodied footprint to dominate. Additionally, the GPU operational carbon typically accounts for the majority of the total operational carbon, while the host system’s embodied carbon footprint dominates the total embodied carbon. This suggests different optimization approaches for operational and embodied carbon respectively: reducing operational carbon requires improving GPU energy efficiency, such as by optimizing workload allocation, minimizing idle power consumption, integrating clean energy sources, and leveraging more energy-efficient GPUs. On the other hand, reducing embodied carbon requires increasing host system utilization through reusing underutilized resources, extending hardware lifetimes, and improving resource sharing to amortize the embodied carbon cost of host components (see Section IV).
Figure 5: Embodied and operational carbon breakdowns of an AI inference system comprising an NVIDIA A100 GPU. We assume a 4-year lifetime running Llama-13B. In data centers and power grids with high degrees of renewable energy (e.g., hyper-scalars), embodied carbon dominates the overall carbon footprint of AI systems.
IV EcoServe Design
Figure 6: Overall EcoServe system diagram. EcoServe takes as input heterogeneous hardware resources, model architectures, and workload traces and optimizes the operational and embodied carbon. EcoServe co-designs four strategies—reusing offline CPUs, rightsizing GPUs, reducing wasted hardware, and asymmetrically recycling hardware—via an ILP formulation to balance performance, energy efficiency, and carbon. The output of EcoServe can be used by external scheduling, and resource provisioning and allocation frameworks to guide system design decisions.
In this section we propose EcoServe, a system that optimizes the holistic carbon footprint of AI systems. Figure 6 shows the overall EcoServe design architecture. EcoServe takes as input LLM models, workload traces and service-level objectives (SLO’s), and hardware platform specifications to reduce carbon footprint while maintain inference performance. EcoServe targets both operational and embodied carbon through four design strategies: reuse, rightsizing, reduce, and recycling. Ecoserve reuses already existing, but unused CPU’s in AI inference systems to process offline/batch workloads, increasing effective throughout and lower resource demand, thus lowering embodied carbon. EcoServe then rightsizes which GPUs across a heterogeneous set (e.g., L4, A6000, A100, H100) to provision for online and offline inference. Based on underutilized host compute, memory, and storage resources and their high embodied overheads (see Section III-B), we reduce wasted resources to further lower carbon footprint. Finally, we propose asymmetrically extending hardware lifetimes and recycling host processing systems and GPUs to balance operational and embodied emissions of AI inference systems. The four design strategies represent a combination of software-only (reuse), resource provisioning and scheduling (right-sizing and recycling), and hardware design (reduce) techniques.
Figure 7: Roofline models of an Intel Sapphire Rapids CPU with 112 cores compared to an NVIDIA A100. We overlay different operators in the decode (solid and dash-dot line) and prompt computation phase (dashed line) for Llama-3-8b with the maximum batch-size that CPU/GPU can hold with context length 2048. At large batch sizes, the GPU is capacity bound ($\star$) and the roofline shows reusing CPUs ($\star$) for offline decode computation is practical.
IV-A Reusing CPUs for Offline Inference
@@
Recall, Section III-B shows host processing systems incur high embodied carbon overheads in inference systems. To maximize the utility of these resources, in this section we propose leveraging host systems, which typically go underutilized, to process offline inference. During auto-regressive inference, CPUs are responsible for the following operations: (1) waiting for incoming requests and processing HTTP and GRPC operations, (2) input tokenization, (3) prefix matching and request ordering, (4) allocating tensors, memory, and meta data for batches, and (5) inference post-processing including de-tokenization, speculative decoding, and beam search processing. Profiling several models (i.e., opt125m, Metallama-3-8B, Mistral and Mixtral) and request rates (2-32 requests per second from ShareGPT [13]) on vLLM (v0.6.2) and SGLang [75], we find average CPU utilization is around 6% of a 16-core processor (e.g., on average only a single CPU core is used). Thus, CPU processing is not on the critical path of recent LLM inference schedulers [68].
Identifying opportunities for using CPUs. Given the underutilization, we must identify which operations can be offloaded from GPUs to CPUs to maximize the benefit. To guide this decision, Figure 7 shows the memory bandwidth and compute gap of an Intel Sapphire Rapids 112-core CPU with an A100 40G GPU. Given the gap between memory bandwidth is lower between the two systems, we find low-arithmetic intensity operations such as attention scoring and decoding phase operators more suitable for CPU inference. Furthermore, the gap between GPU and CPU performance is decreased when running large batch operations given the GPU’s lower capacity. Figure 7 shows the maximal batch-size of 16 and 512 for GPU and CPU running at context length 2048 in FP16.
Figure 8: CPU offline inference requires carefully balancing parallelism degree (PD) and arithmetic intensity of Linear operators (AI, FLOPS/Byte). The run-time schedule can co-design the tile and workload slice dimensions with architecture decisions to optimize for offline performance.
Optimizing CPU performance via better parallelization. To optimize CPU offline inference performance, we must carefully balance the degree of thread-level parallelism exploited per operation, as well as the operational arithmetic intensity, compute requirements, and memory bandwidth requirements. This optimal configuration depends on hardware (e.g., AMX and AVX efficiency, memory system), model, batch sizes, and sequence lengths of input queries. The optimal partitioning strategy involves carefully configuring the thread-level parallelism and tensor tiling dimensions to maximize performance. Figure 8 shows the impact of tile-size on arithmetic intensity (AI) and parallelism degree (PD) based on an analytical performance model assuming a hidden dimension size of 4096. In LLM’s QKV Projection, given an (M, K) and (K, N) GEMM, we focus on parallelizing and tiling across $N$ and $K$ Furthermore, attention operators also constitute a large part of latency, especially for long context workloads, where CPUs exhibit more benefit [11, 65, 16, 15, 32]. Parallelization along the keys/values sequence length dimension, in addition to the batch dimension, maximizes memory bandwidth utilization across all CPU sockets.
@@
Following the roofline analysis, while prompt computation is not desirable on the CPU, we can move the KV cache in a layered pipelined fashion from GPUs to CPUs to perform decoding. Optimizing these parallelism dimensions we find can improve CPU performance by up to 3.67$\times$ compared to a Llama.cpp baseline [25]; across an array of model sizes and sequence lengths, we observe a 1.4$\times$ improvement in performance (Figure 14). As CPU inference optimization is a nascent field, we expect future efforts leveraging sparsity and other optimizations can further improve CPU offline performance.
Offline inference demand: Given their low utilization, we propose using available CPUs to perform offline inference. Central to exploiting CPUs for offline inference are two key observations: demand for offline versus online inference, and dinstinct SLOs for online and offline inference. Figure 9 illustrates the ratio of demand for offline and online inference for two services (A and B) in a real, production cloud provider over a week (left) and within a day (right). On average, offline inference accounts for 21% and 45% of total demand and infrastructure capacity for Service A and B, respectively. Note, at some points in the week, the offline demand peaks at around 27% and 55% for Service A and B respectively; shaving resource demands at this peak can reduce resource requirements. By leveraging available, underutilized CPUs and DRAM in existing systems, we reduce infrastructure capacity needed for offline inference given their relaxed SLOs.
Since offline workloads do not have stringent SLO, it is beneficial to run large batches on CPU and reduce the GPU provisioned. In practise, we can also offload offline demands back to GPU once the online GPU servers are idle for high carbon intensity regions, since GPUs are more energy-efficient. This runtime scheduler decision can be jointly designed with workload partitioner.
@@
Figure 9: Breakdown of online and offline demand for two LLM services (A and B) running in a production, cloud data center over a week (left) and during a day (right). Offline demand account for a significant portion of overall Generative AI serving load, which can be processed by less performant hardware (e.g., CPUs).
Load-aware reuse: Figure 10 shows the server capacity (overprovisioned by 30%) needed for offline infrastructure to serve Llama-3-8B models under time varying demands based on production loads for Service A shown in Fig 9. We show the impact of using CPUs to either process offline inferences during peak demand (red) or continuously at all times (blue). We assume the provision reallocation happens every 4 hour. EcoServe can use CPUs from nodes where GPUs are processing both offline and online inferences to gain significant infrastructure capacity. Reusing CPUs during peak demand, we find can reduce offline capacity demands by nearly 1.32$\times$, saving significant embodied carbon at-scale. The capacity reduction shown is conservative, assuming fixed batch sizes for CPUs and GPUs (for shorter context length, larger models); maximizing batch size on CPUs can further reduce capacity by up to 45%.
IV-B Right-sizing GPU Provisioning
In this section we investigate provisioning heterogeneous GPU and scaling based on varying resource demands to optimize not just performance and efficiency, but also carbon footprint. Fundamentally, recent work has shown different model architectures, sequence lengths, and LLM phases (i.e., prompting and decoding) have different compute, memory bandwidth, and memory capacity requirements [62, 53, 52, 76]. Furthermore, modern GPUs have diverse hardware characteristics that impact performance, performance-per-watt, and performance-per-carbon tradeoffs.

### v2

@@ 3.1.2. Memory and Storage
Recent efforts from industry demonstrate memory and storage dominate the embodied carbon of Microsoft’s data center hardware (greensku, ). To quantify the memory and storage embodied carbon, we consider both memory technologies and capacities. For storage, we estimate SSD emissions based on the Dell R740 life cycle analysis (dellr740, ), yielding an average estimate of 0.110 kg CO2e per GB capacity; this is a conservative estimate compared to estimates of 0.160 kg CO2e per GB from academic surveys (dirty_ssd, ). For memory, we extend existing ACT and SCARIF models by scaling wafer-scale DRAM data for different technologies from TechInsights (TechInsights, ; techinsights_micron_dram_2024, ; skhynixHynixNewsroom, ) by publicly available DRAM bit-densities per technology from memory vendors (e.g., SK Hynix, Micron, Samsung) as shown in Figure 3. The trends show newer technology nodes, given higher bit-densities, exhibit lower embodied carbon per GB capacity. For the purposes of this work we estimate the DDR4/LPDDR5, GDDR6, and HBM2, HBM3e embodied carbon footprint as 0.29, 0.36, 0.28, 0.24 kg CO2e per GB respectively.
@@
Figure 4. Trends in embodied carbon, power, and cloud cost for different generations of GPUs. As GPU performance increases (left to right), power consumption, cost and embodied carbon exhibit distinct trends, ACT only accounts for around 20% in the blue SoC component  (aws-gpu, ; azure-gpu, ; lambdalabsCloudDeep, ).
3.1.3. Mainboard Printed Wiring Board (PCB)
To estimate the embodied emissions of the mainboard printed wiring board (PWB) we use the Dell R740 LCA (dellr740, ). Given the PWB area of the Dell R740, 1925 square centimeters (176 kg CO2), is significantly higher than modern server-class host systems and GPU’s, we scale the PWB area for specific systems. Thus, we estimate the embodied carbon of PWB to be: Total PCB $\text{Total PCB CF}=0.048\times Area_{PCB}$.
3.2. Characterizing Total Carbon
@@
Figure 4 analyzes the embodied carbon estimates of various NVIDIA GPUs, breaking it down into key components such as memory, printed circuit board (PCB), power delivery network (PDN), cooling systems, and system-on-chip (SoC). The breakdown aligns with findings from the recent TPU embodied carbon modeling paper (schneider2025life, ). Figure 4 illustrates these components using stacked bars to represent embodied carbon, with the thermal design power (TDP) plotted on the right axis. We observe a clear trend: with newer GPU generations, computational performance, memory bandwidth, and energy efficiency improve. However, the embodied carbon also rises with increasing computational performance, advanced thermal solutions required, and improved memory capacity. For instance, compared to an NVIDIA H100, and NVIDIA L4 incurs 3$\times$ lower embodied carbon (see Section 4.1.2).
@@
Figure 5. Embodied carbon breakdown of full inference systems available in cloud offerings from Azure and LambdaLabs (azure-gpu, ; lambdalabsCloudDeep, ), varying the number and type of GPU. Host-processing systems account for over half of the embodied carbon in AI systems, owing largely to memory, storage, and mainboard overheads.
Figure 5 shows the embodied carbon breakown of end-to-end inference systems including both GPUs and host systems. The analysis highlights that the motherboard, DRAM, and storage constitute a significant portion of the embodied carbon impact. However, for AI inference workloads, host systems are frequently underutilized compared to GPUs, leading to a disproportionate contribution of their embodied carbon footprint relative to their actual usage during operations. This observation motivates exploring strategies such as resource sharing on the host (e.g., model multiplexing), improving utilization through CPU-based inference, extending the lifetime of host system components, and reducing the memory subsystem’s footprint (Section 4.1.3).
@@
Figure 6 compares embodied and operational carbon per second across datacenters with different power sources. The host and GPU embodied carbon remain fixed, while operational carbon varies with energy source carbon intensity. In high-carbon regions (Europe, US-East), GPU operational carbon dominates. As cleaner energy is adopted, operational emissions drop, making embodied carbon more significant. GPU power accounts for most operational carbon, while the host system dominates embodied carbon. Optimizing operational carbon requires improving GPU energy efficiency, minimizing idle power consumption, dynamic workload allocation, and clean energy integration. Reducing embodied carbon involves increasing host utilization, extending hardware lifetimes, and improving resource sharing to amortize the embodied carbon cost of host components (see Section 4.1.4).
@@
Observation 4: CPU is underutilized in AI inference systems. During auto-regressive inference, CPUs are responsible for the following operations: (1) waiting for incoming requests and processing HTTP and GRPC operations, (2) input tokenization, (3) prefix matching and request ordering, (4) allocating tensors, memory, and metadata for batches, and (5) inference post-processing including de-tokenization, speculative decoding, and beam search processing. Profiling several models (i.e., opt125m, Metallama-3-8B, Mistral and Mixtral) and request rates (2-32 requests per second from ShareGPT (sharegpt, )) on vLLM (v0.6.2) and SGLang (zheng2023efficiently, ), we find average CPU utilization is around 6% of a 16-core processor (e.g., on average only a single CPU core is used). Newer optimization like overlapped schedulers are proposed, so CPU processing is not on the critical path of recent LLM inference schedulers (vllm2024perfupdate, ). For multimodality, encoder-based models or agentic systems, CPU can have a higher utilization and thus in itself provide an opportunity to reuse the CPU side of the host system, while recycling or reducing the memory subsystem (See Section 4.1.1).
Figure 6. Embodied and operational carbon breakdowns of an AI inference system comprising an NVIDIA A100 GPU. We assume a 4-year lifetime running Llama-13B. In data centers and power grids with high degrees of renewable energy (e.g., hyper-scalars), embodied carbon dominates the overall carbon footprint of AI systems.
4. EcoServe Design
Figure 7. EcoServe system diagram: An optimization framework that minimizes operational and embodied carbon across hardware resources through 4R co-designed strategies. Outputs inform scheduling and resource allocation decisions.
EcoServe is a carbon-aware framework for serving large language models (LLMs) that jointly optimizes performance, cost, and environmental impact through systematic resource management. The framework operates on four key principles: Reuse, Right Size, Reduce, and Recycle (4R), orchestrating both software and hardware provisioning decisions to minimize the carbon footprint of LLM inference. Figure 7 shows the overall EcoServe design architecture.
EcoServe’s framework ingests hardware specifications, LLM characteristics, and production traces alongside carbon intensity data to make carbon-aware scheduling decisions. The output such as number of GPUs, CPU cores for each workloads are directly usable for commodity resource scheduler or autoscaler.
4.1. Design Principles
EcoServe implements four strategies for carbon-efficient inference:
  • Reuse (Software Runtime): Leverages idle CPU capacity in existing AI inference systems for offline/batch workloads, increasing throughput while reducing resource demand and embodied carbon.
  • Right Size (Software Provisioning): Optimizes GPU provisioning across heterogeneous hardware (L4, A6000, A100, H100, etc.) for both online and offline inference, matching resources to workload characteristics.
  • Reduce (Hardware SKU Design): Minimizes waste by eliminating underutilized host compute, memory, and storage resources that contribute to high embodied carbon overhead.
  • Recycle (Hardware Provisioning): Extends hardware lifetimes asymmetrically across host processing systems and GPUs to balance operational and embodied emissions.
These strategies provide a comprehensive solution across software runtime optimization, software provisioning, and hardware provisioning to reduce the carbon footprint of AI inference systems. Below we detail on each strategy, then present how the output of the strategy interacts with the components in real systems.
4.1.1. Reusing CPUs for Offline Inference
Section 3.2 shows that host processing systems incur high embodied carbon overheads in AI inference infrastructure. To maximize the utility of these resources, EcoServe opportunistically reuses host CPUs for offline inference, reducing the reliance on high-power GPUs.
Identifying opportunities for using CPUs. Due to the underutilization of host resources, we must carefully determine which operations or inference phases can be offloaded from GPUs to CPUs to maximize carbon efficiency. Note that each offload incurs additional energy costs for data movement, necessitating a selective approach to CPU reuse.
To guide this decision, Figure 8 presents a roofline model analysis (williams2009roofline, ; yuan2024llm, ; nerscRooflinePerformance, ) comparing the computational intensity and memory bandwidth efficiency of an Intel Sapphire Rapids 112-core CPU and an NVIDIA A100 40GB GPU. Given the relatively lower memory bandwidth gap between the CPU and GPU, we find that low-arithmetic-intensity operations, such as attention scoring and KV cache decoding, are well-suited for CPU execution. Additionally, for large-batch operations, GPU throughput becomes capacity-bound, making CPU reuse a practical alternative for offline inference. Figure 8 highlights the maximum batch sizes feasible for GPUs (16) and CPUs (512) at a context length of 2048 in FP16.
Figure 8. Roofline models of an Intel Sapphire Rapids CPU with 112 cores compared to an NVIDIA A100. We overlay different operators in the decode (solid and dash-dot line) and prompt computation phase (dashed line) for Llama-3-8B with the maximum batch size that CPU/GPU can handle at a context length of 2048. At large batch sizes, the GPU is capacity bound ($\star$) while the CPU ($\star$) can efficiently process offline decoding workloads.
Optimizing CPU performance via parallelization. To maximize the efficiency of CPU-based inference, EcoServe optimally configures thread-level parallelism and tensor tiling to balance:
  • Arithmetic intensity (FLOPs/Byte),
  • Compute vs. memory bandwidth constraints,
  • Model size, batch size, and sequence length.
The optimal configuration depends on the CPU microarchitecture (e.g., AMX and AVX efficiency), number format (FP6, BF6, INT8, etc), and workload characteristics. As illustrated in Figure 9, parallelization along the KV sequence length dimension, in addition to the batch dimension, maximizes memory bandwidth utilization across all CPU cores. This is particularly beneficial for long-context workloads, where attention mechanisms contribute significantly to inference latency (flashdecoding2023, ; flashinfer, ; dao2022flashattention, ; dao2023flashattention, ; hong2023flashdecoding++, ).
@@ 4.1.1. Reusing CPUs for Offline Inference
Following our roofline analysis, while prompt computation remains GPU-favorable, EcoServe implements layered pipelining to move KV cache transfers from GPUs to CPUs for decoding. By optimizing these parallelism dimensions, we achieve a 3.67$\times$ speedup compared to a baseline Llama.cpp implementation (llama_cpp, ), with an average 1.4$\times$ improvement across model sizes and sequence lengths (Figure 18). As CPU inference remains a nascent field, future advancements leveraging sparsity and weight-sharing techniques are expected to further enhance CPU-based offline inference (song2023powerinfer, ; zhang2024h2o, ; Liu2023DejaVu, ).
Figure 9. CPU offline inference requires carefully balancing parallelism degree (PD) and arithmetic intensity of Linear operators (AI, FLOPS/Byte). The run-time scheduler co-designs tile and workload slice dimensions with architecture decisions to optimize offline performance.
Adapting to fluctuating offline inference demand. Given their inherently relaxed SLOs, offline workloads offer an opportunity to leverage CPU capacity during periods of low CPU utilization. EcoServe’s reuse strategy is motivated by two key insights:
  1. (1)
    Offline inference demand is significant: Figure 10 illustrates the ratio of online vs. offline demand in a real-world cloud deployment over a week (left) and a single day (right). On average, offline workloads account for 21% and 45% of total infrastructure capacity in Services A and B, respectively.
  2. (2)
    Offline workloads exhibit time-varying peaks: At peak periods, offline demand reaches 27% and 55%, respectively. By scheduling offline inference to underutilized CPU nodes, EcoServe reduces peak GPU resource requirements, thereby lowering total embodied carbon.
Since offline workloads lack stringent latency constraints, EcoServe batches large offline requests on CPUs, thereby reducing the need for high-GPU provisioning. For datacenters with high-carbon-intensity power grids, EcoServe can dynamically reallocate workloads back to GPUs when energy efficiency becomes a priority.
@@
Figure 10. Breakdown of online and offline demand for two LLM services (A and B) running in a production cloud data center over a week (left) and during a day (right). Offline demand accounts for a significant portion of total serving capacity, creating opportunities for CPU reuse in offline inference.
Load-aware reuse. To quantify the impact of CPU reuse, Figure 11 presents the required infrastructure capacity for serving Llama-3-8B inference workloads under varying demand conditions. We analyze two scenarios:
  • Peak-aware reuse (red): CPUs handle offline inference only during peak demand periods.
  • Continuous reuse (blue): CPUs process offline inference at all times.
Assuming resource reallocation occurs every 4 hours, EcoServe’s load-aware CPU reuse reduces offline GPU provisioning needs by 1.32$\times$ at peak demand, contributing to substantial embodied carbon savings. This estimate is conservative, assuming fixed batch sizes for CPU and GPU inference. By further increasing CPU batch sizes, offline capacity reductions of up to 45% are achievable.
Figure 11. Impact of CPU reuse strategies on infrastructure capacity. Red and blue curves represent peak-only and continuous reuse, respectively. EcoServe dynamically reallocates workloads to CPUs, reducing offline GPU capacity demands by up to 1.32$\times$.
4.1.2. Right-sizing GPU Provisioning
While CPU reuse amortized embodied carbon, efficient GPU provisioning is equally critical. EcoServe optimizes GPU allocation by considering three key factors: energy grid carbon intensity, workload characteristics, and hardware diversity. In regions with predominantly renewable energy, CPU reuse is prioritized; in fossil fuel-dependent regions, workloads may shift to more energy-efficient GPUs to minimize overall emissions.
LLM has distinct phases (prompting, decoding) and workload has various characteristics (architecture, sequence length)–this impose different compute and memory requirements. By dynamically matching these requirements to diverse GPU capabilities, EcoServe reduces both operational and embodied carbon while maintaining performance targets.